OpenStudy (anonymous):

solve the problem. use 3.14 for pie, and round your answer to the nearest tenth. the foundation for a cylindrical flower bed is a cylinder 17 feet in diameter and 5 feet high. how much concrete is needed to pour the foundation? a. 533.8 cu. ft b. 2268.7 cu. ft c. 1134.3 cu. ft d. 4537.3 cu. ft

OpenStudy (anonymous):

someone plllzzzz help!!! WILL GIVE MEDAL!!!!

jimthompson5910 (jim_thompson5910):

Volume of Cylinder V = pi*r^2*h In this case, pi = 3.14 approximately r = 8.5 (divide the diameter 17 in half) h = 5

OpenStudy (anonymous):

so do i just multiply those 3 numbers??

jimthompson5910 (jim_thompson5910):

yes and the notation "r^2" means you square the radius example: 5^2 = 5*5 = 25

OpenStudy (anonymous):

do i square it first before multiplying it all??

jimthompson5910 (jim_thompson5910):

yes exponents come before multiplication PEMDAS P = parenthesis E = exponents M = multiplication D = division A = addition S = subtraction

OpenStudy (anonymous):

its 1134 so c right??!! thats what i got

jimthompson5910 (jim_thompson5910):

you should get 1,134.325

jimthompson5910 (jim_thompson5910):

1,134.325 rounded to the nearest tenth is 1,134.3

jimthompson5910 (jim_thompson5910):

looks good

OpenStudy (anonymous):

yea i know but i rounded

OpenStudy (anonymous):

ty!!

jimthompson5910 (jim_thompson5910):

np